A snowmobile is originally at the point with position vector 28.5 m at 95.0° counterclockwise from the x axis, moving with velocity 4.69 m/s at 40.0°. It moves with constant acceleration 1.85 m/s2 at 200°. After 5.00 s have elapsed, find the following. (Express your answers in vector form.)
1
Expert's answer
2019-09-30T10:06:29-0400

After 5 seconds velocity is given by

"\overset{\to}{v}=\overset{\to}{u}\ +\ 5\overset{\to}{a}"

Talking about components

"v_x=u_x\ +\ 5a_x\\v_y=u_y\ +\ 5a_y"

"v_x=4.69\cos(40\degree)+5\times1.85\cos(200\degree)\\v_y=4.69\sin(40\degree)+5\times 1.85\sin(200\degree)"

"v_x=-5.098\\v_y=-0.149\\|v|=\sqrt{v_x^2+v_y^2}=5.1\ m\ sec^{-1}"

"\tan(\theta)=\frac{0.149}{5.098}=0.0292\\\theta=1.67\\since \ both\ component\ are\ negative\ so angle \ will\ be\ 180+1.67=181.67\degree" So velocity after 5 seconds is elapsed will be

"5.1\ m\ sec^{-1}" at an angle of "181.67\degree"

Or in term of vector component

"\overset{\to}{v}=-5.098\hat{i}-0.149\hat{j}"